Is the point (-5,1) a solution of -3y+2x<-9 (The < has a line under it.)
Answer by stanbon(75887)

Is the point (-5,1) a solution of -3y+2x<-9 (The < has a line under it.)
------------
Substitute for x and for y to see if you get a true statement:
--------------
-3*1 + 2*-5 <= -9
-3 -10 <= -9
-13 <= -9
True
-----------------
Therefore (-5,1) is an element of the solution set.
===================================================
Cheers,
Stan H.